What mass of silver chloride can be produced from 1.32 L of a 0.273 M solution of silver nitrate?
1
Expert's answer
2019-10-02T08:02:56-0400

Number of moles of silver nitrate"=Vol \times Molarity=1.32 \times 0.273=0.36036"

Reaction to displace nitrate to chloride:

"AgNO_3+Cl^-\to AgCl+NO_3^-"

so ,Number of moles of silver chloride formed"=" Number of moles of silver nitrate used=0.36036

Mass of silver chloride formed"=Molar \ mass\times Number\ of \ moles=0.36036 \times 143.32=51.647 gm"